OpenStudy (anonymous):

A movie theater sells popcorn in small, medium, large and jumbo sizes. The customers of the first show purchase 4 small, 20 medium, 40 large, and 16 jumbo containers of popcorn. Find the probability shown below. p(large container)

OpenStudy (anonymous):

\[P(\large container) = 40/80=1/2 =50%\] 40 is number of choices we have for large container and 80 is out of all sizes. got it?